y=4x
x+y=5 solve each system of linear eguations using subsitution

Answer by fractalier(6550)

Starting from
y=4x
x+y=5
we substitute the first into the second and get
x + 4x = 5
5x = 5
x = 1
Then
y = 4(1) = 4